Whenever a photon is emitted by hydrogen in Balmer series,it is followed by another photon Lyman series.why?

Grade:12

1 Answers

Gaurav
9 years ago
Hello Student
Balmer series corresponds to the wavelengths in the visible spectrum and Lyman corresponds to ones in the Ultraviolet.
the equations for each are
Lyman series : 1/λ = R(1 - 1/n2) n = 2,3,4...
Balmer series : 1/λ = R(1/4 - 1/n2) n = 3,4,5...